Please help... thanks in advance!

Suppose a simple random sample 0 size n-75 is obtained rom a population whose size s 一20,0 0 and whose population propo tion with a specified characten cs p Oo om iete parts hrough be。 (a) Describe the sampling distribution of p. Choose the phrase that best describes the shape of the sampling distribution below. o A. Not normal because ns 005N and np(1-p) 2 10. B. Approximately normal because n 0.05N and np(1-p) < 10 O C. Approximately normal because n0.05N and np(1 -p)210 O D. Not normal because n 0.05N and np(1-p) < 10
Answer #1

Correct answer: option (C) Approximately normal because n < = 0.5 N and npq >=10

since

n = 75 < 0.05N = 0.05*20000 = 1000

and npq = 75*0.6*0.4 = 18 >= 10